# Generalization of DNNs:

### v1

@@ A.2 Model setups
Appendix B AccNet Generalization Bounds
@@
Consider an accordion net of depth $L$ and widths $d_{L},\dots,d_{0}$, with corresponding set of functions =$\mathcal{F}=\{f_{L:1}:\left\|f_{\ell}\right\|_{F_{1}}\leq R_{\ell},\text{Lip}(f_{\ell})\leq\rho_{\ell}\}$ with input space $\Omega=B(0,r)$. For any $\rho$-Lipschitz loss function $\ell(x,f(x))$ with $\left|\ell(x,y)\right|\leq c_{0}$, we know that with probability $1-\delta$ over the sampling of the training set $X$ from the distribution $\pi$, we have for all $f\in\mathcal{F}$
$\mathcal{L}(f)-\tilde{\mathcal{L}}_{N}(f)$ $\leq C\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}\frac{\log N}{\sqrt{N}}(1+o(1))+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}.$
@@
The strategy is: (1) prove a covering number bound on $\mathcal{F}$ (2) use it to obtain a Rademacher complexity bound, (3) use the Rademacher complexity to bound the generalization error.
@@
Let us now consider a sequence $\epsilon_{k}=2^{-k}$ for $k=0,\dots,K$ and define $\tilde{v}_{\ell,i}^{(k)},\tilde{w}_{\ell,i}^{(k)}$ to be the $\epsilon_{k}$-covers of $\bar{v}_{\ell,i},\bar{w}_{\ell,i}$, furthermore we may choose $\tilde{v}_{\ell,i}^{(0)}=\tilde{w}_{\ell,i}^{(0)}=0$ since every unit vector is within a $\epsilon_{0}=1$ distance of the origin. We will now show that on can approximate $f_{\theta}$ by approximating each of the $f_{\ell}$ by functions of the form
@@
And for the induction step, we condition on the layers $f_{\ell-1:1}$
@@
(2) Our goal now is to use chaining / Dudley’s theorem to bound the Rademacher complexity $R(\mathcal{F}(X))$ evaluated on a set $X$ of size $N$ (e.g. Lemma 27.4 in [42]) of our set:
Lemma 7.
Let $c=\max_{f\in\mathcal{F}}\frac{1}{\sqrt{N}}\left\|f(X)\right\|$, then for any integer $M>0$,
$R(\mathcal{F}(X))\leq c2^{-M}+\frac{6c}{\sqrt{N}}\sum_{k=1}^{M}2^{-k}\sqrt{\log\mathcal{N}(\mathcal{F},c2^{-k})}.$
To apply it to our setting, first note that for all $x\in B(0,r)$, $\left\|f_{L:1}(x)\right\|\leq\rho_{L:1}r$ so that $c=\max_{f\in\mathcal{F}}\frac{1}{\sqrt{N}}\left\|f(X)\right\|\leq\rho_{L:1}r$, we then have
$R(\mathcal{F}(X))$ $\leq c2^{-M}+\frac{6c}{\sqrt{N}}\sum_{k=1}^{M}2^{-k}12s_{0}\sqrt{e}\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}c^{-1}2^{k}(1+o(1))$
$=c2^{-M}+\frac{72}{\sqrt{N}}Ms_{0}\sqrt{e}\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}(1+o(1)).$
Taking $M=\left\lceil-\log_{2}\left(\frac{72}{\sqrt{N}}s_{0}\sqrt{e}\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}\right)\right\rceil$, we obtain
$R(\mathcal{F}(X))$ $\leq\frac{72}{\sqrt{N}}Ms_{0}\sqrt{e}\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}(1+M(1+o(1)))$
$\leq\frac{144}{\sqrt{N}}Ms_{0}\sqrt{e}\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}\left\lceil-\log_{2}\left(\frac{72}{\sqrt{N}}s_{0}\sqrt{e}\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}\right)\right\rceil(1+o(1))$
$\leq C\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}\frac{\log N}{\sqrt{N}}(1+o(1)).$
(3) For any $\rho$-Lipschitz loss function $\ell(x,f(x))$ with $\left|\ell(x,y)\right|\leq c_{0}$, we know that with probability $1-\delta$ over the sampling of the training set $X$ from the distribution $\pi$, we have for all $f\in\mathcal{F}$
$\mathbb{E}_{x\sim\pi}\left[\ell(x,f(x))\right]-\frac{1}{N}\sum_{i=1}^{N}\ell(x_{i},f(x_{i}))$ $\leq 2\mathbb{E}_{X^{\prime}}\left[R(\ell\circ\mathcal{F}(X^{\prime}))\right]+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}$
$\leq 2C\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}\frac{\log N}{\sqrt{N}}(1+o(1))+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}.$
Appendix C Composition of Sobolev Balls

### v2

@@ A.2 Model setups
Figure 4: Comparison of our bounds to a number of previous bounds on the composition of two functions dataset across different differentiabilities of $g$ and $h$. For simplicity, we drop the constant prefactors in each bound, since they are mostly an artifact of the proofs. We see that our bound (Jacot24 Jacobian) obtains strictly better than previous ones, even when using upper bounding the Lipschitz constants by the operator norms (Jacot24 OP). Since we cannot compute the Lipschitz constant, we approximate them by a max over 100 random points (taking more points does not significantly change the final bound). The lines appear flat because most of the variation is between different bounds rather than as $N$ increases, but by zooming in, we can see a downward trend, albeit a slow and noisy one. The case $(2,2)$ lies in the regime where both $g$ and $h$ are hard, in which case our theory predicts that the operator norm bound should have a worse scaling exponent than the Lipschitz based bound. The experiments appear to match our prediction, because the Lipschitz-constant based bound seems to be the only one that is decreasing in $N$, while the operator norm bound, as well as all previous bounds, are increasing (the fact that uniform generalization bounds can be increasing in $N$ has been observed in previous work (Nagarajan & Kolter, 2019)). Though these trends are interesting, there is clearly a lot of noise in the resulting curves, so it is difficult to confidently conclude anything from these experiments.
Appendix B Covering Numbers
We will use a covering number argument to prove generalization bounds. Before we start, we must define a few notions.
The $\epsilon$-covering number $\mathcal{N}_{2}(\mathcal{F};\epsilon)$ of set $\mathcal{F}$ of functions $f:\Omega\to\mathbb{R}^{d}$, is the smallest integer $n$ such that for any distribution $\pi$ supported on $\Omega$ there is a covering $\tilde{\mathcal{F}}$, i.e. a finite set with no more than $n$ elements, such that for all $f\in\mathcal{F}$ there is a cover $\tilde{f}\in\tilde{\mathcal{F}}$ with
$\left\|f-\tilde{f}\right\|_{\pi}=\sqrt{\mathbb{E}_{x\sim\pi}\left\|f(x)-\tilde{f}(x)\right\|^{2}}\leq\epsilon.$
We will also sometimes say that $\tilde{f}$ covers $f$ in such case.
The covering number allows us to bound the difference between the errors of two datasets uniformly over the whole functions set $\mathcal{F}$ with high-probability:
Theorem 6.
Let $\mathcal{F}$ be a space of functions such that $\left\|g\right\|_{\infty}\leq B$ for all $g\in\mathcal{F}$. Given a true function $f^{*}$ such that $\left\|f\right\|_{\infty}\leq B$, we have that with probability at least $1-p$ over the sampling of two datasets $S,S^{\prime}$ from a distribution $\pi$, one has for all $g\in\mathcal{F}$:
$\left\|g-f^{*}\right\|_{S^{\prime}}-\left\|g-f^{*}\right\|_{S}\leq 2\min_{\epsilon}\epsilon+B\sqrt{2\frac{\log\mathcal{N}(\mathcal{F},\epsilon)-\log p}{N}}.$
Proof.
We start with a Chernoff bound:
$\mathbb{P}_{S,S^{\prime}}\left[\sup_{g\in\mathcal{F}}\left\|g-f^{*}\right\|_{S^{\prime}}-\left\|g-f^{*}\right\|_{S}\geq c\right]\leq\mathbb{E}_{S,S^{\prime}}\left[\exp\left(t\sup_{g\in\mathcal{F}}\left\|g-f^{*}\right\|_{S^{\prime}}-\left\|g-f^{*}\right\|_{S}\right)\right]e^{-tc}$
Defining $M(t)=\mathbb{E}_{S,S^{\prime}}\left[\exp\left(t\sup_{g\in\mathcal{F}}\left\|g-f^{*}\right\|_{S^{\prime}}-\left\|g-f^{*}\right\|_{S}\right)\right]$, this can be reformulated as saying that with probability $1-p$, we have that for all $g\in\mathcal{F}$
$\left\|g-f^{*}\right\|_{S^{\prime}}-\left\|g-f^{*}\right\|_{S}<\min_{t}\frac{\log M(t)-\log p}{t}.$
Let us now bound $M(t)$. For any pair of datasets $S,S^{\prime}$, we consider the $\epsilon$-covering $\tilde{\mathcal{F}}$ of $\mathcal{F}$ w.r.t. to the measure $\frac{S+S^{\prime}}{2}$ which is the empirical measure corresponding to the dataset of size $2N$ made up of the union of $S$ and $S^{\prime}$. Thus for any function $g\in\mathcal{F}$, there is a $\tilde{g}\in\tilde{\mathcal{F}}$ such that $\left\|g-\tilde{g}\right\|_{\frac{S+S^{\prime}}{2}}\leq\epsilon$ and |~$\left|\tilde{\mathcal{F}}\right|\leq\mathcal{N}(\mathcal{F},\epsilon)$.
We can now bound
$M(t)$ $=\mathbb{E}_{S,S^{\prime}}\left[\exp\left(t\sup_{g\in\mathcal{F}}\left\|g-f^{*}\right\|_{S^{\prime}}-\left\|g-f^{*}\right\|_{S}\right)\right]$
$\leq\mathbb{E}_{S,S^{\prime}}\left[\exp\left(t\sup_{g\in\mathcal{F}}\left\|\tilde{g}-g\right\|_{S}+\left\|\tilde{g}-g\right\|_{S^{\prime}}+\left\|\tilde{g}-f^{*}\right\|_{S^{\prime}}-\left\|\tilde{g}-f^{*}\right\|_{S}\right)\right]$
$\leq e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\exp\left(t\sup_{g\in\mathcal{F}}\left\|\tilde{g}-f^{*}\right\|_{S^{\prime}}-\left\|\tilde{g}-f^{*}\right\|_{S}\right)\right]$
$=e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\sup_{g\in\mathcal{F}}\exp\left(t\frac{\sum_{i=1}^{N}\left\|\tilde{g}(x_{i})-f^{*}(x_{i})\right\|^{2}-\left\|\tilde{g}(x_{i}^{\prime})-f^{*}(x_{i}^{\prime})\right\|^{2}}{N\left(\left\|\tilde{g}-f^{*}\right\|_{S^{\prime}}+\left\|\tilde{g}-f^{*}\right\|_{S}\right)}\right)\right]$
$\leq e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\sup_{g\in\mathcal{F}}\exp\left(t\frac{\sum_{i=1}^{N}\left\|\tilde{g}(x_{i})-f^{*}(x_{i})\right\|^{2}-\left\|\tilde{g}(x_{i}^{\prime})-f^{*}(x_{i}^{\prime})\right\|^{2}}{\sqrt{2}N\left\|\tilde{g}-f^{*}\right\|_{\frac{S+S^{\prime}}{2}}}\right)\right]$
where we used the fact that $\left\|h\right\|_{S}+\left\|h\right\|_{S^{\prime}}\leq\sqrt{2\left\|h\right\|_{S}^{2}+2\left\|h\right\|_{S^{\prime}}^{2}}=2\left\|h\right\|_{\frac{S+S^{\prime}}{2}}\leq 2\epsilon$ and conversely $\left\|h\right\|_{S}+\left\|h\right\|_{S^{\prime}}\geq\sqrt{\left\|h\right\|^{2}_{S}+\left\|h\right\|^{2}_{S^{\prime}}}=\sqrt{2}\left\|h\right\|_{\frac{S+S^{\prime}}{2}}$. We may now introduce i.i.d. Rademacher random variables $\sigma_{i}$ equal to $1$ and $-1$ with prob. $\frac{1}{2}$:
$M(t)$ $\leq e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\mathbb{E}_{\sigma}\left[\sup_{\tilde{g}\in\tilde{\mathcal{F}}}\exp\left(t\frac{\sum_{i=1}^{N}\sigma_{i}\left(\left\|\tilde{g}(x_{i})-f^{*}(x_{i})\right\|^{2}-\left\|\tilde{g}(x_{i}^{\prime})-f^{*}(x_{i}^{\prime})\right\|^{2}\right)}{\sqrt{2}N\left\|\tilde{g}-f^{*}\right\|_{\frac{S+S^{\prime}}{2}}}\right)\right]\right]$
$\leq e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\sum_{\tilde{g}\in\tilde{\mathcal{F}}}\prod_{i=1}^{N}\mathbb{E}_{\sigma_{i}}\left[\exp\left(\frac{t\sigma_{i}\left(\left\|\tilde{g}(x_{i})-f^{*}(x_{i})\right\|^{2}-\left\|\tilde{g}(x_{i}^{\prime})-f^{*}(x_{i}^{\prime})\right\|^{2}\right)}{\sqrt{2}N\left\|\tilde{g}-f^{*}\right\|_{\frac{S+S^{\prime}}{2}}}\right)\right]\right].$
To understand why adding the Rademacher random variables $\sigma_{i}$ does not change the expectation, one can think of these random variables as randomly switching the $i$-th datapoints $x_{i}$ from $S$ and $x^{\prime}_{i}$ from $S^{\prime}$ with each other (not that these switches leave the denominator unchanged too). Sampling two datasets $S$ and $S^{\prime}$ followed by random switches is equivalent to sampling without the switches.
We now use Hoeffdings Lemma, which tells us that for any random var $X$ with $\mathbb{E}X=0$ and $X\in[-b,b]$ a.s. one has $\mathbb{E}\exp tX\leq\exp\frac{t^{2}b^{2}}{2}$ to obtain
$M(t)$ $\leq e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\sum_{\tilde{g}\in\tilde{\mathcal{F}}}\exp\left(\frac{t^{2}\sum_{i=1}^{N}\left(\left\|\tilde{g}(x_{i})-f^{*}(x_{i})\right\|^{2}-\left\|\tilde{g}(x^{\prime}_{i})-f^{*}(x^{\prime}_{i})\right\|^{2}\right)^{2}}{2N^{2}\left\|\tilde{g}-f^{*}\right\|_{\frac{S+S^{\prime}}{2}}^{2}}\right)\right]$
$\leq e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\sum_{\tilde{g}\in\tilde{\mathcal{F}}}\exp\left(\frac{t^{2}B^{2}\sum_{i=1}^{N}\left\|\tilde{g}(x_{i})-f^{*}(x_{i})\right\|^{2}+\left\|\tilde{g}(x^{\prime}_{i})-f^{*}(x^{\prime}_{i})\right\|^{2}}{N^{2}\left\|\tilde{g}-f^{*}\right\|_{\frac{S+S^{\prime}}{2}}^{2}}\right)\right]$
$=e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\sum_{\tilde{g}\in\tilde{\mathcal{F}}}\exp\left(\frac{t^{2}B^{2}\left(\left\|\tilde{g}-f^{*}\right\|_{S}^{2}+\left\|\tilde{g}-f^{*}\right\|_{S^{\prime}}^{2}\right)}{N\left\|\tilde{g}-f^{*}\right\|_{\frac{S+S^{\prime}}{2}}^{2}}\right)\right]$
$\leq e^{2t\epsilon}\mathbb{E}_{S,S^{\prime}}\left[\sum_{\tilde{g}\in\tilde{\mathcal{F}}}\exp\left(\frac{t^{2}2B^{2}}{N}\right)\right]$
$\leq\mathcal{N}(\mathcal{F},\epsilon)\exp\left(2t\epsilon+\frac{t^{2}2B^{2}}{N}\right).$
This bound on $M(t)$ then implies that with probability at least $1-p$ over the sampling of $S,S^{\prime}$ for all $g\in\mathcal{F}$
$\left\|g-f^{*}\right\|_{S^{\prime}}-\left\|g-f^{*}\right\|_{S}$ $\leq\min_{t}\frac{\log M(t)-\log p}{t}$
$\leq\min_{t,\epsilon}2\epsilon+\frac{2B^{2}t}{N}+\frac{\log\mathcal{N}(\mathcal{F},\epsilon)-\log p}{t}$
$=2\min_{\epsilon}\epsilon+B\sqrt{2\frac{\log\mathcal{N}(\mathcal{F},\epsilon)-\log p}{N}}.$
We can then use the above to bound the generalization error uniformly:
Theorem 7.
Under the same setup as Theorem 6, we have that with prob. $1-p$ for all $g\in\mathcal{F}$
$\left\|g-f^{*}\right\|_{\pi}-\left\|g-f^{*}\right\|_{S}\leq\min_{\epsilon}2\epsilon+B\sqrt{8\frac{\log\mathcal{N}(\mathcal{F},\epsilon)}{N}}+cB\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}},$
for $c=\sqrt{2}(3+\sqrt{5})$.
Proof.
Our goal is to approximate the population error $\left\|f^{*}-g\right\|_{\pi}^{2}$ by the test error $\left\|f^{*}-g\right\|_{S^{\prime}}^{2}$ for a sample $S^{\prime}$ independent of $g$, and this test error is itself close to the train error thanks to Theorem 6.
We use Bennett’s inequality which relies on the fact that the function $h(x)=\left\|f^{*}(x)-g(x)\right\|^{2}$ is bounded by $4B^{2}$:
$\mathbb{P}\left[\left\|f^{*}-g\right\|_{\pi}^{2}-\left\|f^{*}-g\right\|_{S^{\prime}}^{2}\geq a\right]$ $=\mathbb{P}\left[\sum_{i=1}^{N}h(x_{i})-N\mathbb{E}h(x)\geq Na\right]$
$\leq\exp\left(-\frac{N\mathbb{E}_{\pi}\left[(h(x)-\mathbb{E}h(x))^{2}\right]}{(4B^{2})^{2}}\phi\left(\frac{4B^{2}a}{\mathbb{E}_{\pi}\left[(h(x)-\mathbb{E}h(x))^{2}\right]}\right)\right)$
for $\phi(u)=(1+u)\log(1+u)-u$.
Lemma 8 gives us the bound $\phi^{-1}(v)\leq\sqrt{2v}+2v$, which implies that with probability $p$
$\left\|f^{*}-g\right\|_{\pi}^{2}-\left\|f^{*}-g\right\|_{S^{\prime}}^{2}$ $\leq\frac{\mathbb{E}_{\pi}\left[(h(x)-\mathbb{E}h(x))^{2}\right]}{4B^{2}}\left(\sqrt{\frac{-2(4B^{2})^{2}\log p}{N\mathbb{E}_{\pi}\left[(h(x)-\mathbb{E}h(x))^{2}\right]}}+\frac{-2(4B^{2})^{2}\log p}{N\mathbb{E}_{\pi}\left[(h(x)-\mathbb{E}h(x))^{2}\right]}\right)$
$=\sqrt{2\mathbb{E}_{\pi}\left[(h(x)-\mathbb{E}h(x))^{2}\right]\frac{-\log p}{N}}+8B^{2}\frac{-\log p}{N}.$
We now use the facts that $\mathbb{E}_{\pi}\left[(h(x)-\mathbb{E}h(x))^{2}\right]\leq\mathbb{E}_{\pi}\left[h(x)^{2}\right]\leq 4B^{2}\mathbb{E}_{\pi}\left[h(x)\right]=4B^{2}\left\|g-f^{*}\right\|_{\pi}^{2}$ to obtain
$\left\|f^{*}-g\right\|_{\pi}^{2}-\left\|f^{*}-g\right\|_{S^{\prime}}^{2}\leq B\left\|g-f^{*}\right\|_{\pi}\sqrt{8\frac{-\log p}{N}}+8B^{2}\frac{-\log p}{N}.$
Reordering this is implies
$\left\|f^{*}-g\right\|_{\pi}^{2}-\left\|f^{*}-g\right\|_{\pi}\sqrt{8B^{2}\frac{-\log p}{N}}-\left\|f^{*}-g\right\|_{S^{\prime}}^{2}-8B^{2}\frac{-\log p}{N}\leq 0$
which by the quadratic formula yields the inequality
$\left\|f^{*}-g\right\|_{\pi}$ $\leq\sqrt{2B^{2}\frac{-\log p}{N}}+\sqrt{2B^{2}\frac{-\log p}{N}+\left\|f^{*}-g\right\|_{S^{\prime}}^{2}+8B^{2}\frac{-\log p}{N}}$
$\leq\left\|f^{*}-g\right\|_{S^{\prime}}+(\sqrt{2}+\sqrt{10})B\sqrt{\frac{-\log p}{N}}$
This implies that with probability $1-2p$ we have that for all $g\in\mathcal{F}$,
$\left\|g-f^{*}\right\|_{\pi}-\left\|g-f^{*}\right\|_{S}$ $\leq\left\|f^{*}-g\right\|_{S^{\prime}}-\left\|g-f^{*}\right\|_{S}+(\sqrt{2}+\sqrt{10})B\sqrt{\frac{-\log p}{N}}$
$\leq\min_{\epsilon}2\epsilon+2B\sqrt{2\frac{\log\mathcal{N}(\mathcal{F},\epsilon)-\log p}{N}}+(\sqrt{2}+\sqrt{10})B\sqrt{\frac{-\log p}{N}}$
$\leq\min_{\epsilon}2\epsilon+B\sqrt{8\frac{\log\mathcal{N}(\mathcal{F},\epsilon)}{N}}+\sqrt{2}(3+\sqrt{5})B\sqrt{\frac{-\log p}{N}}.$
Let us now prove the lemma we used in the proof:
Lemma 8.
The function $\phi(u)=(1+u)\log(1+u)-u$ satisfies
$\phi(u)\leq\frac{1}{1+u}\frac{u^{2}}{2}.$
It is also invertible over the non-negative reals and its inverse satisfies
$\phi^{-1}(v)\leq\sqrt{2v}+2v.$
Proof.
To obtain the first inequality, we compute the derivatives of $\phi$:
$\phi^{\prime}(u)$ $=\log(1+u)+1-1=\log(1+u)$
$\phi^{\prime\prime}(u)$ $=\frac{1}{1+u}.$
which implies that
[…]
Solving for a non-negative $u$ in
$v=\frac{1}{1+u}\frac{u^{2}}{2}$
yields
$2v=\frac{1}{u^{-2}+u^{-1}}$
$\implies$ $u^{-2}+u^{-1}-\frac{1}{2v}=0$
$\implies$ $u^{-1}=\frac{-1+\sqrt{1+\frac{2}{v}}}{2}$
$\implies$ $u=\frac{2\sqrt{v}}{\sqrt{2+v}-\sqrt{v}}=\sqrt{v}(\sqrt{2+v}+\sqrt{v})$
which implies that
$\phi^{-1}(v)\leq\sqrt{v}(\sqrt{2+v}+\sqrt{v})\leq\sqrt{2v}+2v.$
Example 1.
As an example consider a function space $\mathcal{F}$ whose log covering number (this is also called the metric entropy) grows polynomially:
$\log\mathcal{N}_{2}(\mathcal{F},\epsilon)\leq c\epsilon^{-r}.$
Therefore we get that
$\left\|f^{*}-f\right\|_{\pi}-\left\|f^{*}-f\right\|_{S}\leq\min_{\epsilon}2\epsilon+\sqrt{8}B\frac{\sqrt{c}}{\sqrt{\epsilon^{r}N}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}.$
Choosing $\epsilon=\left(2B^{2}\frac{c}{N}\right)^{\frac{1}{2+r}}$ (chosen so that the two terms depending on $\epsilon$ match), we obtain
$\left\|f^{*}-f\right\|_{\pi}-\left\|f^{*}-f\right\|_{S}\leq 2\left(2B^{2}\frac{c}{N}\right)^{\frac{1}{2+r}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}},$
which is of order $N^{-\frac{1}{2+r}}$.
Assuming that the train error is zero (or sufficiently small, i.e. $O(N^{-\frac{1}{2+r}})$), we get that
$\left\|f^{*}-f\right\|_{\pi}^{2}=O(N^{-\frac{2}{2+r}}).$
Appendix C AccNet Generalization Bounds
@@
Consider an accordion net of depth $L$ and widths $d_{L},\dots,d_{0}$, with corresponding set of functions =$\mathcal{F}=\{\sigma_{L}\circ f_{L:1}:\left\|f_{\ell}\right\|_{F_{1}}\leq R_{\ell},\text{Lip}(f_{\ell})\leq\rho_{\ell}\}$ with input space $\Omega=B(0,r)$ and with a final nonlinearity $\sigma_{L}$ that is $1$-Lipschitz and uniformly bounded by $B$. Then with probability $1-p$ over the sampling of the training set $X$ from the distribution $\pi$, we have for all $f\in\mathcal{F}$
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $\leq c_{0}\sqrt{B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}N^{-\frac{1}{2}}}(1+o(1))+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N},}$
for $c_{0}\approx 14.6$ and $c_{1}\approx 7.4$. Thus if ~$\sqrt{\tilde{\mathcal{L}}_{N}(f)}\leq c_{0}\sqrt{B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}N^{-\frac{1}{2}}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}$ we have
$\mathcal{L}(f)$ $\leq 8c_{0}^{2}B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}N^{-\frac{1}{2}}(1+o(1))+8c_{1}^{2}B^{2}\frac{-\log\nicefrac{{p}}{{2}}}{N}.$
@@
The strategy is: (1) prove a covering number bound on $\mathcal{F}$, (2) use it to bound the generalization error.
@@
Let us now consider a sequence $\epsilon_{k}=2^{-k}$ for $k=0,\dots,K$ and define $\tilde{v}_{\ell,i}^{(k)},\tilde{w}_{\ell,i}^{(k)}$ to be the $\epsilon_{k}$-covers of $\bar{v}_{\ell,i},\bar{w}_{\ell,i}$, furthermore we may choose $\tilde{v}_{\ell,i}^{(0)}=\tilde{w}_{\ell,i}^{(0)}=0$ since every unit vector is within a $\epsilon_{0}=1$ distance of the origin. We will now show that one can approximate $f_{\theta}$ by approximating each of the $f_{\ell}$ by functions of the form
@@
And for the induction step, we condition on the first $\ell-1$ layers
@@
(2) We now apply Theorem 7 along the lines of Example 1 with $r=2$ and
$c=144s_{0}^{2}e\rho_{L:1}^{2}r^{2}\left(\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}\right)^{2}\frac{1}{\delta^{2}},$
to obtain
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $\leq 2\sqrt{24Bs_{0}\sqrt{e}\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}}N^{-\frac{1}{4}}(1+o(1))$
$+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}$
$=c_{0}\sqrt{B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}}N^{-\frac{1}{4}}(1+o(1))+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N},}$
for $c_{0}=2\sqrt{24s_{0}\sqrt{e}}\approx 14.6$ and $c_{1}=\sqrt{2}(3+\sqrt{5})\approx 7.4$.
Therefore if ~$\sqrt{\tilde{\mathcal{L}}_{N}(f)}\leq c_{0}\sqrt{B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}}N^{-\frac{1}{4}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}$, we have
$\mathcal{L}(f)$ $\leq\left(2c_{0}\sqrt{B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}}N^{-\frac{1}{4}}(1+o(1))+2c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}\right)^{2}$
$\leq 8c_{0}^{2}B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}N^{-\frac{1}{2}}(1+o(1))+8c_{1}^{2}B^{2}\frac{-\log\nicefrac{{p}}{{2}}}{N}.$
Appendix D Composition of Sobolev Balls
Let us first prove a simple generalization bound for Sobolev balls: